# Discovering environments with XRM

### v1

@@ 4.1 Twin setup, holding-out of data
We start by initializing two twin label predictors $f^{a}$ and $f^{b}$. Without loss of generality, let these predictors return softmax probability vectors over the $n_{\text{classes}}$ classes in the training data. We split our training dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$ in two random halves. Formally, we construct a pair of training assignment vectors with entries $m^{a}_{i}\sim\text{Bernoulli}(\frac{1}{2})$ and $m^{b}_{i}=1-m^{a}_{i}$, for all $i=1,\ldots,n$. For predictor $f^{a}$, examples with $m^{a}_{i}=1$ are “held-in” and examples with $m^{a}_{i}=0$ are “held-out”; similarly for $f^{b}$. Therefore, we will train predictor $f^{a}$ on training examples where $m^{a}_{i}=1$, and similarly for predictor $f^{b}$. Before learning starts, we calibrate the softmax temperature of the twins via Platt scaling (Guo et al., 2017). See appendix D for implementation details.
@@
By virtue of this arrangement, we we may now estimate the generalization difficulty of any example by looking at the prediction of the twin that held-out such point. This contrasts prior methods for environment discovery, which consume the entire training data, and may therefore conflate generalization and memorization. (Relatedly, Feldman and Zhang (2020) proposes a similar “error when holding-out” construction as a measure of memorization.) Particularly to our interests, if a point is misclassified when held-out, we see this as evidence of such example belonging to the minority group. As a last remark, we recommend choosing the twins to inhabit the same function class as the downstream phase-2 DG predictor. This allows discovering environments tailored to the point of view of the chosen learning machine.
4.2 Twin training, flipping labels
As figure 1 shows, the test worst-group-accuracy of an ERM baseline on Waterbirds is $62\%{}$. This suggests that, if using ERM to train our twins, each would be able to correctly classify roughly one half of the minority examples. If using these machines to discover environments based on prediction errors, we would dilute the spurious correlation evenly across the two discovered environments. Consequently, it would be difficult for a phase-2 DG algorithm to tell apart between invariant and spurious patterns. Albeit counter-intuitive, we would like to hinder the learning process of our twins, such that they increasingly rely on spurious correlation. In the best possible case, the twins would correctly classify all majority examples and mistake all minority examples, resulting in zero worst-group accuracy.
To this end, we propose to steer away our twins from becoming empirical risk minimizers as follows. Let $p^{\text{out}}_{i}=m^{b}_{i}f^{a}(x_{i})+m^{a}_{i}f^{b}(x_{i})$ be the held-out softmax prediction for example $(x_{i},y_{i})$. Also, let $y^{\text{out}}_{i}=\arg\max_{j}p^{\text{out}}_{i,j}$ be the held-out predicted class label, equal to the index of the maximum held-out softmax prediction. Then, at each iteration during the training of the twins,
flip $\text{flip }y_{i}\text{ into }y_{i}^{\text{out}},\text{ with probability }(p^{\text{out}}_{y^{\text{out}}_{i}}-1/n_{\text{classes}})\cdot n_{\text{classes}}/(n_{\text{classes}}-1),$ (3)
and let each twin take a gradient step to minimize their held-in class-balanced—according to the moving targets—cross-entropy loss.
The overarching intuition is that the label flipping equation 3 implements an “echo chamber” reinforcing the twins to rely on spurious correlation. Label flipping happens more often for confident held-out mistakes and early in training. These are two footprints of spurious correlations, since these are often easier and faster to capture. (In the context of neural networks, this is often referred to as a “simplicity bias” (Shah et al., 2020).) Overall, the purpose of equation 3 is to transform the labels of the training data such that they do not longer represent the original classes, but spurious bias. Finally, the adjustment of equation 3 in terms of $n_{\text{classes}}$ ensures low flip probabilities at initialization, where most mistakes are due to weight randomness, and not due to spurious correlation.
4.3 Twin model selection, counting label flips
Before discovering environments, we must commit to a pair of twin predictors. Since these have their own hyper-parameters, XRM would be incomplete without a phase-1 model selection criterion (Gulrajani and Lopez-Paz, 2020). We propose to select the twin hyper-parameters showing a maximum number of label flips (3) at the last iteration, and across the training data. To reiterate, by “counting flips” we simply compare the vector of current labels with the vector of original labels—therefore, we do not accumulate counts of double or multiple flips per label. To understand why, recall that each label flip signifies one example that is confidently misclassified when held-out. Therefore, each label flip is evidence about reliance on spurious correlation, which consequently brings us closer to a clear-cut identification of the minority group.
4.4 Environment discovery, using cross-mistake formula
Having committed to a pair of twins, we are ready to discover environments for all of our training and validation examples. In particular, we use a simple “cross-mistake” formula to annotate any example $(x,y)$ with the binary environment
$e(x,y)=\llbracket(y\notin\mathrm{argmax}_{j}f^{a}(x)_{j})\,\lor\,(y\notin\mathrm{argmax}_{j}f^{b}(x)_{j})\rrbracket.$ (4)
[…]
We are now ready to train the phase-2 DG algorithm of our choice on top of the training data with environments discovered with XRM. When doing so, we can perform phase-2 DG model selection by maximizing worst-group-accuracy on the validation data with environments discovered by XRM.
5 Experiments
Our experimental protocol has three moving pieces: datasets, phase-2 domain generalization algorithms, and the source of environment annotations.
Datasets
We consider six standard datasets from the SubpopBench suite (Yang et al., 2023). These are the four image datasets Waterbirds (Wah et al., 2011), CelebA (Liu et al., 2015), MetaShift (Liang and Zou, 2022), and ImageNetBG (Xiao et al., 2020); and the two natural language datasets MultiNLI (Williams et al., 2017) and CivilComments (Borkan et al., 2019). We invite the reader to Appendix B.1 from Yang et al. (2023) for a detailed description of these. For CelebA, predictors map pixel intensities into a binary “blonde/not-blonde” label. No individual face characteristics, landmarks, keypoints, facial mapping, metadata, or any other information was used to train our CelebA predictors. We also conduct experiments on ColorMNIST (Arjovsky et al., 2019), but keep a strict protocol. More specifically, we set both training and validation data to contain two environments, with $0.8$ and $0.9$ label-color correlation, while the test environment shows $0.1$ label-color correlation. This contrasts Arjovsky et al. (2019), who used the test environment for model selection.
Phase-2 DG algorithms
We consider ERM, group distributionally robust optimization (Sagawa et al., 2019, GroupDRO), group re-weighting (Japkowicz, 2000, RWG), and group sub-sampling (Idrissi et al., 2022, SUBG). When group information is available, we tune hyper-parameters and early-stopping by maximizing worst-group-accuracy. Otherwise, we tune for worst-class-accuracy. Following standard praxis, image datasets employ a pretrained ResNet-50, while text datasets use a pretrained BERT. For more details, see appendix A.
Environment annotations
For each combination of dataset and phase-2 DG algorithm, we compare group annotations from different sources. None denotes no group annotations. Human denotes ground-truth annotations, as originally provided in the datasets, and inducing oracle performance. XRM denotes group annotations from the environments discovered by our proposed method. In some experiments we compare XRM to other environment discovery methods, these being learning from failure (Nam et al., 2020, LfF), environment inference for invariant learning (Creager et al., 2020, EIIL), just train twice (Zheran Liu et al., 2021, JTT), automatic feature re-weighting (Qiu et al., 2023, AFR), and learning to split (Bao and Barzilay, 2022, LS).
Metrics
Regardless of how training and validation groups are discovered, we always report test worst-group-accuracy over the human group annotations provided by each dataset. The tables hereby presented show averages over ten random seeds. For tables with error bars, see appendix C.
5.1 XRM versus human annotations
Section 5.1 shows that XRM enables oracle-like worst-group-accuracy across datasets. The performance gains are remarkable in the challenging ColorMNIST dataset, where XRM perfectly identifies digits appearing in minority colors, discovering a pair of environments conducive of stronger generalization than the ones originally proposed by humans. For the commonly-reported quartet of Waterbirds, CelebA, MultiNLI, and CivilComments, human annotations induce an average oracle worst-group-accuracy $80.6\%{}$, while XRM environments endow a super-human performance of $80.9\%{}$.
Table 1: Worst-group-accuracies for all datasets, algorithms, and environment annotations, averages over ten random seeds. XRM discovers environments conducive of oracle-like DG performance.
@@ 5.1 XRM versus human annotations
Waterbirds 70.4 76.1 75.3 71.7 88.0 86.1 74.8 87.0 84.5 73.0 86.7 76.3
CelebA 62.7 71.9 67.6 68.8 89.1 89.8 70.7 89.5 88.0 68.5 87.1 87.5
MultiNLI 54.8 65.3 65.8 69.7 76.1 74.3 69.3 71.1 73.3 53.7 72.8 71.3
CivilComments 55.1 59.7 61.4 59.3 69.3 71.3 54.0 65.8 73.4 58.7 64.0 72.9
ColorMNIST 10.1 10.1 26.9 10.0 10.1 70.5 10.1 10.2 70.2 10.1 10.0 70.8
MetaShift 70.8 67.7 71.2 70.8 73.7 71.2 66.5 70.8 69.7 70.0 73.5 69.2
ImagenetBG 75.6 76.9 76.9 73.0 76.1 75.9 76.9 76.5 77.0 75.0 75.0 76.4
Average 57.1 61.1 63.6 60.5 68.9 77.0 60.3 67.3 76.6 58.4 67.0 74.9
5.2 XRM versus other methods for environment discovery
@@
Section 5.2 shows the worst-group-accuracy of GroupDRO when built on top of environments as discovered by different methods. As seen in the previous subsection, XRM achieves $80.4\%{}$, nearly matching oracle performance. The second best method with no access to environment information, JTT, drops to $58.9\%{}$. The best method accessing a validation set with human environment annotations, AFR, lags far from XRM, with $78\%{}$. The computational burden to complete the results from LS was prohibitive. For example, one run of LS for Waterbirds, the smallest dataset, took 20 hours. An XRM run for this same dataset, on the same 32GB Volta GPU, takes 10 minutes. As detailed in appendix B, we found that by varying some of the fixed hyper-parameters in the official LS repository, performance on Waterbirds varied by as much as ±$\pm 7\%{}$ in worst-group accuracy.
@@
ERM 86.1 76.1 93.5 71.9 78.6 65.3 82.9 59.7 85.3 68.3
GroupDRO 92.6 88.0 93.3 89.1 82.0 76.1 81.4 69.3 87.3 80.6
${}^{\dagger}$ 97.3 72.6 95.6 47.2 82.4 67.9 83.1 69.5 89.6 64.3
${}^{\dagger}$ 91.2 78.0 85.1 77.2 80.8 70.2 68.2 50.3 81.3 68.9
${}^{\dagger}$ 96.9 78.7 89.5 77.8 79.4 70.0 90.5 67.0 89.1 73.4
${}^{\dagger}$ 93.3 86.7 88.0 81.1 78.6 72.6 83.3 64.3 85.8 76.2
${}^{\dagger}$ 94.4 90.4 91.3 82.0 81.4 73.4 89.8 68.7 89.2 78.6
ERM 85.3 70.4 94.5 62.7 77.9 54.8 80.9 55.1 84.6 60.8
${}^{\dagger}$ 86.6 75.0 81.1 53.0 71.4 57.3 69.1 42.2 77.1 56.9
${}^{\dagger}$ 90.8 64.5 95.7 41.7 80.3 64.7 — — — —
${}^{\dagger}$ 88.9 71.2 95.9 48.3 81.4 65.1 79.0 51.0 86.3 58.9
${}^{\dagger}$ 91.2 86.1 87.2 83.3 78.7 72.1 — — — —
XRM 90.6 86.1 91.8 91.8 78.3 74.3 79.9 71.3 85.2 80.9
5.3 Some visualizations
@@
Figure 2: XRM on the Waterbirds problem, concerning   waterbirds in water,   waterbirds in land,   landbirds in water,   landbirds in land. The first panel shows that “percentage of XRM label flipped at convergence” is a strong indicator of “worst-group-accuracy in phase-2”, making flips a good criterion to select twin hyper-parameters. The two middle panels show the signed margin of the twins on each ground-truth group. From each of these class-dependent plots, XRM discovers two environments: one for points in the “mistake-free” white area, and one for points in the “cross-mistake” gray areas. Notably, XRM is able to allocate the two smallest groups    to dedicated environments. The fourth panel shows that label flipping happens almost exclusively for the two smallest groups, and stabilizes as training progresses.
Figure 3: Randomly selected images of CIFAR-10 from groups identified by XRM. The twin networks show interesting patterns in their mistakes. Notably, well-classified examples are prototypical, such as planes against blue backgrounds and deers against a green backgrounds.
@@
Figure 2 explores some of the behaviors of XRM on the Waterbirds dataset. In particular, the left panel justifies the use of “percentage of label flipped at convergence” as a phase-1 model selection criterion for XRM, as it correlates strongly with downstream phase-2 worst-group-accuracy. The two middle panels showcase the clear separation of the minority group “landbirds/water” by XRM, as no landbirds in land are in the cross-mistake area. The right panel shows that label flipping happens almost exclusively for minority groups, and converges alongside XRM training. This provides XRM with a degree of stability, removing the need for intricate early-stopping criteria.
@@
Figure 3 applies XRM to the CIFAR-10 dataset (Krizhevsky et al., 2009). While CIFAR-10 does not contain environment annotations, the discovered environments by XRM for the “plane” and “deer” classes reveal one interesting spurious correlations, namely background color. As a final remark, we ablated the need for (i) holding-out data, and (ii) performing label flipping, finding that both components are essential to the performance of XRM.

### v2

@@ 4.1 Twin Setup, Holding-out of Data
We start by initializing two twin classifiers $f^{a}$ and $f^{b}$. Without loss of generality, let these classifiers return softmax probability vectors over the $n_{\text{classes}}$ classes in the training data. We split our training dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$ in two random halves. Formally, we construct a pair of training assignment vectors with entries $m^{a}_{i}\sim\text{Bernoulli}(\frac{1}{2})$ and $m^{b}_{i}=1-m^{a}_{i}$, for all $i=1,\ldots,n$. For classifier $f^{a}$, examples with $m^{a}_{i}=1$ are “held-in” and examples with $m^{a}_{i}=0$ are “held-out”; similarly for $f^{b}$. Therefore, we will train classifier $f^{a}$ on training examples where $m^{a}_{i}=1$, and similarly for classifier $f^{b}$. See Appendix C for implementation details.
@@
By virtue of this arrangement, we may now estimate the generalization difficulty of any example by looking at the prediction of the twin that held-out such point. This contrasts prior methods, which consume the entire training data, and may therefore conflate generalization and memorization. Here, however, if a point is misclassified when held-out, we see this as evidence of such example belonging to the minority group. Feldman & Zhang (2020) proposes a similar “error when holding-out” construction as a measure of memorization. In the context of label-noise robustness, CrossSplit (Kim et al., 2023) also employs a similar approach, in which, confident held-out mistakes are indicators of a model’s memorization of a noisy label.
Table 1: Worst-group-accuracies, averaged from ten runs across datasets and algorithms, show XRM achieving oracle-level performance. When group labels are not available, class labels substitute them. Additionally, while ERM does not use group labels, it can still benefit from validation group labels for hyperparameter tuning, resulting in improved performance.
@@
Waterbirds
66.4 66.4 66.4 67.3 86.5 88.1 66.8 87.3 82.0 62.1 87.1 77.7
CelebA 54.3 55.1 58.6 68.4 88.3 89.1 64.7 82.6 81.1 65.9 83.9 81.1
MultiNLI 67.9 72.0 69.1 68.6 73.4 72.1 66.9 71.3 70.2 69.7 52.4 72.0
CivilComments 67.2 74.0 64.7 66.6 73.8 72.2 67.2 73.4 72.2 65.4 71.1 44.6
ColorMNIST 10.0 10.1 11.3 9.9 10.1 69.7 10.0 10.3 69.4 10.0 9.9 65.2
MetaShift 64.2 69.8 71.7 72.5 80.3 77.5 72.3 76.3 74.2 69.7 77.0 77.6
ImagenetBG 78.4 78.4 79.0 78.3 78.2 78.2 78.3 79.0 78.8 79.5 78.8 77.2
Average 58.3 60.8 60.1 61.6 70.1 78.1 60.9 68.6 75.4 60.3 65.8 70.8
4.2 Twin Training, Flipping Labels
As Figure 1 shows, the test worst-group-accuracy of an ERM baseline on Waterbirds is $62\%{}$. This suggests that, if using ERM to train our twins, each would be able to correctly classify roughly one half of the minority examples. If using these machines to discover environments based on prediction errors, we would dilute the spurious correlation evenly across the two discovered environments. Consequently, it would be difficult for an OOD generalization algorithm to tell apart between invariant and spurious patterns. Albeit counter-intuitive, we would like to hinder the learning process of our twins, such that they increasingly rely on spurious correlation. In the best possible case, the twins would correctly classify all majority examples and misclassify all minority examples, resulting in zero worst-group accuracy.
To this end, we propose to steer away our twins from becoming empirical risk minimizers as follows. Let $p^{\text{out}}_{i}=m^{b}_{i}f^{a}(x_{i})+m^{a}_{i}f^{b}(x_{i})$ be the held-out softmax prediction for example $(x_{i},y_{i})$. Also, let $y^{\text{out}}_{i}=\arg\max_{j}p^{\text{out}}_{i,j}$ be the held-out predicted class label. Then, at each iteration during the training of the twins, flip $y_{i}$ into $y_{i}^{\text{out}}$, with probability,
$(p^{\text{out}}_{y^{\text{out}}_{i}}-1/n_{\text{classes}})\cdot n_{\text{classes}}/(n_{\text{classes}}-1),$ (2)
and let each network to minimize their held-in cross-entropy loss—according to the moving targets.
The overarching intuition is that the label flipping Equation 2 implements an “echo chamber” reinforcing the twins to rely on spurious correlation. Label flipping happens more often for confident held-out mistakes and early in training. These are two footprints of spurious correlations, since these are often easier and faster to capture. (In the context of neural networks, this is often referred to as a “simplicity bias” (Shah et al., 2020b; Pezeshki et al., 2021).) Overall, the purpose of Equation 2 is to transform the labels of the training data such that they no longer represent the original classes, but spurious bias. Finally, the adjustment of Equation 2 in terms of $n_{\text{classes}}$ ensures low flip probabilities at initialization, where mistakes are random, and not due to spurious correlation. We note that the “echo chamber” effect aligns the twin networks and that is crucially different from methods that use multiple networks to either disagree with or diversify spurious features (Nam et al., 2020; Cha et al., 2021; Rame et al., 2022; Wortsman et al., 2022; Lee et al., 2023; Pagliardini et al., 2023; Lin et al., 2023; Eastwood et al., 2023).
4.3 Twin Model Selection, Counting Label Flips
Before discovering environments, we must commit to a pair of twin classifiers. Each of the twin networks own hyper-parameters, XRM would be incomplete without a model selection criterion (Gulrajani & Lopez-Paz, 2020). We propose to select the twin hyper-parameters showing a maximum number of label flips at the last iteration, and across the training data. To reiterate, by “counting flips” we simply compare the vector of current labels with the vector of original labels—therefore, we do not accumulate counts of double or multiple flips per label. To understand why, recall that each label flip signifies one example that is confidently misclassified when held-out. Therefore, each label flip is evidence about reliance on spurious correlation, which consequently brings us closer to a clear-cut identification of the minority group.
4.4 The Cross-Mistake Formula
Having committed to a pair of twins, we are ready to discover environments for all of our training and validation examples. In particular, we use a simple “cross-mistake” formula to annotate any example $(x,y)$ with the binary annotation, $e(x,y)=$
$\llbracket(y\notin\mathrm{argmax}_{j}f^{a}(x)_{j})\,\lor\,(y\notin\mathrm{argmax}_{j}f^{b}(x)_{j})\rrbracket,$ (3)
[…]
We are now ready to train the OOD generalization algorithm of our choice on top of the training data with environments discovered with XRM.
5 Experiments
This section presents a series of experiments to showcase the effectiveness of XRM on two well-known benchmarks. Additional experiments are also conducted to identify scenarios where XRM excels, as well as scenarios where it fails to discover relevant environments.
5.1 Sub-population Shift Benchmarks
For sub-population shift tasks, we experiment with seven datasets and four algorithms detailed in Appendix B. We compare results with 3 sources of environment annotations:
  • None: no env. annotations—class labels are used instead,
  • Human: original human-annotated environments,
  • XRM: inferred environments discovered by our method.
Metrics
Regardless of how training and validation environments are discovered, we always report test worst-group-accuracy over the human environment annotations provided by each dataset. The tables hereby presented show averages over ten random seeds. For results with error bars, see Section B.1.
XRM vs. human annotations
Section 4.1 shows that XRM enables oracle-like worst-group-accuracy across datasets. The performance gains are remarkable in the challenging ColorMNIST dataset, where XRM perfectly identifies digits appearing in minority colors, discovering a pair of environments conducive of stronger generalization than the ones originally proposed by humans. For the commonly-reported quartet of Waterbirds, CelebA, MultiNLI, and CivilComments, the average worst-group-accuracy is $67.3\%{}$ when no group annotations are used. When using XRM, the average worst-group-accuracy significantly improves to $80.4\%{}$, closely matching $80.6\%{}$ achieved with human annotations.
XRM vs. other methods
@@ 5.1 Sub-population Shift Benchmarks
Section 5.1 compares the worst-group-accuracy achieved by GroupDRO using XRM-inferred environments against other environment discovery methods. These include learning from failure (Nam et al., 2020, LfF), environment inference for invariant learning (Creager et al., 2020, EIIL), just train twice (Zheran Liu et al., 2021, JTT), correct-n-contrast (Zhang et al., 2022b, CnC), automatic feature re-weighting (Qiu et al., 2023, AFR), and LS (Bao & Barzilay, 2022). As seen in the previous subsection, XRM achieves $80.4\%{}$, nearly matching oracle performance. The second best method with no access to environment information, JTT, drops to $58.9\%{}$. The best method accessing a validation set with human environment annotations, AFR, lags far from XRM, with $78.6\%{}$. The computational burden to complete the results from LS was prohibitive, with more details provided in Appendix B.9. For example, one run of LS for Waterbirds, the smallest dataset, took 20 hours. An XRM run for this same dataset, on the same 32GB Volta GPU, takes 10 minutes.
@@
ERM 83.8 66.4 95.5 55.1 81.6 72.0 84.3 74.0 86.3 66.9
GroupDRO 90.2 86.5 93.1 88.3 80.6 73.4 84.2 73.8 87.0 80.5
ERM 97.3 72.6 95.6 47.2 82.4 67.9 83.1 69.5 89.6 64.3
LfF 91.2 78.0 85.1 77.2 80.8 70.2 68.2 50.3 81.3 68.9
EIIL 96.9 78.7 89.5 77.8 79.4 70.0 90.5 67.0 89.1 73.4
JTT 93.3 86.7 88.0 81.1 78.6 72.6 83.3 64.3 85.8 76.2
CnC 90.9 88.5 89.9 88.8 — — — — — —
AFR 94.4 90.4 91.3 82.0 81.4 73.4 89.8 68.7 89.2 78.6
ERM 83.5 66.4 95.4 54.3 82.1 67.9 81.3 67.2 85.6 63.9
LfF 86.6 75.0 81.1 53.0 71.4 57.3 69.1 42.2 77.1 56.9
EIIL 90.8 64.5 95.7 41.7 80.3 64.7 — — — —
JTT 88.9 71.2 95.9 48.3 81.4 65.1 79.0 51.0 86.3 58.9
LS 91.2 86.1 87.2 83.3 78.7 72.1 — — — —
BAM 91.4 89.1 88.4 80.1 80.3 70.8 88.3 79.3 87.1 79.8
XRM 89.3 88.1 91.4 89.1 75.8 72.1 84.0 72.2 85.1 80.4
5.2 The DomaninBed Benchmark
Section 5.2 presents additional domain generalization results on the DomainBed benchmark (Gulrajani & Lopez-Paz, 2020). Experiments compare three settings: ERM without any environment annotations, the CORAL domain generalization algorithm (Sun & Saenko, 2016) with human-annotated environments, and CORAL with environments discovered by XRM. As a note, CORAL is the best performing single-model (non-ensembling) method in the DomainBed suite. Once again, results suggest that the performance when using XRM-inferred annotations is comparable to that of human-annotated environments. Further details on these experiments are provided in Section B.2 with full table of results in Section B.2.
Table 3: The average and worst test environment accuracies for five datasets in the DomainBed benchmark (Gulrajani & Lopez-Paz, 2020). Three methods are compared: 1) ERM with no environment annotations, 2) CORAL with human-annotated environments, and 3) CORAL with XRM-inferred environments. The model selection is done according to the average accuracy over validation environments.
{NiceTabular}
[…]
Method (annotations) Avg Worst Avg Worst Avg Worst Avg Worst Avg Worst
ERM (None) 77.97 64.85 83.35 72.55 65.47 52.25 47.02 34.60 31.69 9.30
CORAL (Human) 77.87 65.00 84.99 77.70 67.74 53.55 48.51 37.15 41.97 13.25
CORAL (XRM) 77.66 66.15 83.81 77.30 67.01 53.90 49.60 38.00 35.87 11.60
5.3 Further Analytical Experiments
@@
Figure 2: XRM on the Waterbirds problem, concerning   waterbirds in water,   waterbirds in land,   landbirds in water,   landbirds in land. The top-left panel shows that “percentage of XRM label-flips at convergence” is a strong indicator of “worst-group-accuracy of OOD generalization algorithm”, making flips a good criterion to select twin hyper-parameters. The two bottom panels show the signed margin of the twins on each ground-truth group. Each of the bottom plots correspond to one of the classes. Note that a positive margin means correct classification. From each of these class-dependent plots, XRM discovers two environments: one for points in the “mistake-free” white area, and one for points in the “cross-mistake” gray areas. Notably, XRM is able to allocate the two smallest groups    to dedicated environments. Another notable observation is that the two bottom plots appear as straight lines, indicating that the twin networks agree on their predictions. The top-right panel shows that label flipping happens almost exclusively for the two smallest groups, and stabilizes as training progresses.
Figure 3: Randomly selected images from CIFAR-10, identified by XRM. Although CIFAR-10 lacks predefined environment annotations, our method has successfully uncovered intriguing environments. Notably, well-classified examples (when held-out) are prototypical, featuring planes in blue skies and deer on green landscapes. In contrast, misclassified examples (when held-out) are less typical, which means they are correctly classified only when included in the training set.
Label flipping dynamics on Waterbirds
@@
Figure 2 explores some of the behaviors of XRM on the Waterbirds dataset. In particular, the top-left panel justifies the use of “percentage of label-flips at convergence” as a model selection criterion for XRM, as it correlates strongly with downstream worst-group-accuracy. The two bottom panels showcase the clear separation of the minority group “landbirds/water” by XRM, as no landbirds in land are in the cross-mistake area. The top-right panel shows that label flipping happens almost exclusively for minority groups, and converges alongside XRM training. This provides XRM with a degree of stability, removing the need for intricate early-stopping criteria.
Revealing spurious correlations in CIFAR-10 with XRM
@@
Figure 3 applies XRM to the CIFAR-10 dataset (Krizhevsky et al., 2009). While CIFAR-10 does not contain environment annotations, the discovered environments by XRM for the “plane” and “deer” classes reveal one interesting spurious correlation, namely the background color.
As a final remark, we ablated the need for (i) holding-out data, and (ii) performing label flipping, finding that both components are essential to the performance of XRM.
6 When Does XRM Fail?
In the previous section, we showcased the effectiveness of XRM in successfully discovering relevant environments. However, it is important to note that XRM, like other environment discovery methods, relies on certain assumptions. In the absence of these assumptions, these methods, including XRM, may fail to discover relevant environments. In its full generality, the problem of learning invariant predictors in the absence of appropriate environment annotations is indeed impossible (Lin et al., 2022a). To see that, we note that dividing data into invariance-affording environments parallels the problem of discovering the right causal structure in the field of causal inference (Pearl, 2009). To reveal the true causal structure between an invariant feature $X_{\text{inv}}$ and a target $Y$, one must “control-for” a set of variables/environments $E$ that satisfy the following conditional independence statement,
$Y\perp E\mid X_{\text{inv}}.$ (4)
Unfortunately, different causal structures can produce identical observational data (Peters et al., 2017). Because of that, identifying an appropriate $E$ is impossible without admitting extra knowledge about the causal structure behind our data, and XRM is not an exception of such free lunch.
Therefore, we would expect XRM to work well in instances where the inferred environments $E$ satisfy (4), and we should anticipate trouble in those cases where the discovered environments violate (4). On the other hand, as discussed in (Lin et al., 2022a), evaluating (4) requires knowing the invariant feature $X_{\text{inv}}$, which is the variable subject to discovery. This makes (4) difficult to verify in practice when inferring environments $E$, and the best we can do is to offer some canonical examples of successes and failures, that can guide our choices of when to apply XRM.
Here, we exemplify with four different versions of the ColorMNIST dataset (Arjovsky et al., 2019). All four versions instantiate a colored digit classification task, differing on whether the invariant feature is “digit shape” or “digit color”, and which one of these two variables bear the strongest correlation to the target label. Overall, we expect “digit color” to be faster (easier) to learn, leading to generalization issues when “digit shape”—more difficult and slower to learn—is the desired invariant feature.
[…]
Table 4: Four ColoredMNIST versions, where the environment $E$ influences digit shape $S$ and color $C$, forming our input $X=(S,C)$. We depict the causal structure for each dataset version, and the correlation between variables. The invariant feature may be the complex digit shape (CMNIST versions) or the simple digit color (MCOLOR versions), which in turn could bear the strongest or weakest correlation to the target variable—producing four versions of the ColoredMNIST problem. Note that CMNIST-MCOLOR and InverseCMNIST-InverseMCOLOR are indistinguishable from pooled training data alone. At the bottom, test accuracies of ERM, XRM+GroupDRO, and an Oracle which relies solely on the invariant feature.
We show in Table 4 the average-test-accuracy of ERM and XRM followed by GroupDRO for the four versions of the ColoredMNIST dataset. We also show what a hypothetical oracle, relying solely on the invariant feature, would achieve. ERM performs well when the invariant feature is the simplest of the two. XRM performs well when the invariant feature is the most complex of the two. We highlight that the datasets CMNIST and MCOLOR are observationally equivalent from pooled data alone—and a similar remark follows for InverseCMNIST and InverseMCOLOR. This echoes the impossibility results of (Lin et al., 2022a), namely learning invariant predictors in the absence of environment annotations is impossible in its full generality: for instance, based on training data alone, we would never know if we are dealing with InverseCMNIST or InverseMCOLOR, and therefore we are at a loss of whether to apply ERM or XRM. Nevertheless, XRM remains an state-of-the-art solution for those problems were we would like our learning machine to ignore the fastest-to-learn feature, often being a spurious shortcut (Geirhos et al., 2020; Shah et al., 2020a; Pezeshki et al., 2021), in order to focus on more complex patterns with a higher potential for invariance.